### v3

# The Dual of Quantifier Elimination: Boolean Elimination over $\mathbb{C}$ and $\mathbb{R}$

Abstract.
We show that every finite Boolean combination of polynomial equalities and inequalities in $\mathbb{C}^{n}$ admits two uniform normal forms: an $\exists\forall$ form and a $\forall\exists$ form, each using a single polynomial equation. Both forms use only one existentially quantified variable and one universally quantified variable. Optimality results demonstrate that no purely existential or universal normal form is possible over $\mathbb{C}$.
These results extend to sets constructible from entire functions, and to quantifier-free formulas in functional languages over infinite fields of characteristic 0. A corollary shows Zilber’s conjecture on quasiminimality equivalent to its subcase quantifying a single equation.
Over $\mathbb{R}$ and $\mathbb{Q}$, similar results hold, including a singly-quantified $\exists$ form for Boolean combinations of equations and inequations, an $\exists$ form for $\mathbb{R}$ established by prior methods, and other results for order inequalities parallel to the forms over $\mathbb{C}$.
These results provide a dual to classical quantifier elimination: instead of removing quantifiers at the cost of increased Boolean complexity, they remove Boolean structure at the cost of a short, fixed quantifier prefix. The constructions have linear degree bounds and are explicit.
2020 Mathematics Subject Classification:
Primary 03C10; Secondary 03C60, 14P10, 14Q20.
1. Introduction
Quantifier elimination replaces quantified formulas by quantifier-free ones, typically at a cost of more complicated Boolean combinations. This paper studies the complementary direction: eliminating Boolean structure at the cost of introducing a short block of quantifiers.
Our main result here shows that over $\mathbb{C}$, every finite Boolean combination of polynomial equations and inequations can be expressed using one polynomial equation with one existential quantifier and one universal quantifier, with the quantifiers in either order. More precisely:
Theorem 1.1.
[Boolean Elimination over $\mathbb{C}$] Let $\varphi(x)$ be any Boolean combination of polynomial equations and inequations in variables $x\in\mathbb{C}^{n}$. Then there are polynomials
$p,q\in\mathbb{C}[x,a,b]$
such that for all $x\in\mathbb{C}^{n}$,
$\varphi(x)$ $\Longleftrightarrow(\exists a\in\mathbb{C})(\forall b\in\mathbb{C})\,p(x,a,b)=0$
$\varphi(x)$ $\Longleftrightarrow(\forall a\in\mathbb{C})(\exists b\in\mathbb{C})\,q(x,a,b)=0$
For example, if
$\varphi(y,z):=(y=0\wedge z\neq 0)\vee(z=0\wedge y\neq 0)$
then the constructions here show
$\varphi\iff(\exists a)(\forall b)\Big[(1-az)+by\Big]\Big[(1-ay)+bz\Big]=0$
and also
$\varphi\iff(\forall a)(\exists b)\Big[1-b(a-1)(a-2)\Big]\Big[(a-2)yz+(a-1)(1-by)(1-bz)\Big]=0.$
Even more complicated formulas admit normal forms with only $\exists$ and one $\forall$, in either order; having both normal forms provides flexibility for use inside larger formulas. Furthermore the degrees of the resulting polynomials are linearly bounded by the length of the original formula in conjunctive or disjunctive normal form.
The main result may also be generalized complex-geometrically or model-theoretically, with the same proofs.
Theorem 1.2.
[Reformulation in Complex Geometry] Let $\mathcal{A}$ be an algebra of entire functions on $\mathbb{C}^{n}$, closed under polynomial operations. Let $\mathcal{A}[a,b]$ denote the extension of this algebra by polynomials in two complex variables $a$ and $b$. Let $S\subset\mathbb{C}^{n}$ be any constructible set defined over $\mathcal{A}$, i.e. a finite Boolean combination of zero-sets $\{x:f(x)=0\}$ where $f\in\mathcal{A}$. Then $S$ can be represented as a locus of identical vanishing: there is a function $P\in\mathcal{A}[a,b]$ such that
$x\in S\Longleftrightarrow\exists a\in\mathbb{C}\text{ where }P(x,a,b)\text{ vanishes identically for all }b\in\mathbb{C}.$
Dually, $S$ can be represented as a locus of universal solvability: there is a function $Q\in\mathcal{A}[a,b]$ such that
[…]
The applies to several possible algebras, including the algebra of polynomials, the algebra of exponential polynomials, the algebra of entire functions of order $\rho$ (for some $\rho>0$), the algebra of entire functions of finite order, and the algebra of all entire functions.
Theorem 1.3.
[Reformulation in Model Theory] Let $\mathcal{L}$ be a functional language containing $1,+,-,\times$. Let $\mathcal{M}$ be an $\mathcal{L}$-structure satisfying the axioms for an infinite field of characteristic 0. Let $\varphi$ be any quantifier-free formula in $\mathcal{L}$ over the variables $x_{1},\ldots,x_{n}$. Then there are $\mathcal{L}$-terms $p,q$ formed from polynomial combinations of $1,a,b$ and the terms occurring in $\varphi$, such that
$\mathcal{M}\models\varphi(x)$ $\Longleftrightarrow(\exists a)(\forall b)\,p(x_{1},\ldots,x_{n},a,b)=0$
$\mathcal{M}\models\varphi(x)$ $\Longleftrightarrow(\forall a)(\exists b)\,q(x_{1},\ldots,x_{n},a,b)=0$
This applies to several possible languages, including the pure polynomial language, the language adding $\exp$, the language adding $1/\Gamma$, the language adding all univariate entire functions, and the language of all entire functions in any number of variables. In particular, we obtain an equivalence for Zilber’s conjecture on the quasiminimality of $\exp$ (on which see Kirby’s survey for more context).
Corollary 1.1.
Let $\mathcal{L}$ be the language $\{+,-,\times,\exp\}$, and let quantifiers $Q_{i}$ be either $\forall$ or $\exists.$ Then the following are equivalent:
  • For any $n$, any sequence of quantifiers $Q_{i}$, and any quantifier-free formula $\varphi(z,x_{1},\ldots,x_{n})$ in $\mathcal{L}$, the set
    $\{z\in C:Q_{1}x_{1}\ldots Q_{n}x_{n}\,\varphi(z,x_{1},\ldots,x_{n})\}\phantom{=0}$
    is either countable or cocountable.
  • For any $n$, any sequence of quantifiers $Q_{i}$, and any term $t(z,x_{1},\ldots,x_{n})$ in $\mathcal{L}$, the set
    $\{z\in C:Q_{1}x_{1}\ldots Q_{n}x_{n}\,t(z,x_{1},\ldots,x_{n})=0\}$
    is either countable or cocountable.
In other words, Zilber’s conjecture is equivalent to its subcase on the quasiminimality of sets defined by quantifying a single equation, rather than sets defined by quantifying an arbitrary formula.
The rest of the paper is organized as follows: Section 2 establishes notation. Section 3 reviews context. Sections 4 and 5 construct the $\exists\forall$ and $\forall\exists$ polynomials for $\mathbb{C}$ explicitly, and provide bounds on their degrees. Section 6 shows that these quantifier patterns are optimal: there are simple formulas over $\mathbb{C}$ that cannot be expressed by purely existential or purely universal quantification of a single equation.
Section 7 turns to other formally real fields, and shows that any Boolean combination of equations and inequations ($u\neq 0$) over them can be represented with a single $\exists$ and a single equality. Section 8 constructs the $\exists^{d}$ and $\forall\exists$ polynomials for Boolean combinations over $\mathbb{R}$ that may involve inequalities ($u>0$). Sections 9 and 10 explore prior results over $\mathbb{R}$ and forms over $\mathbb{Q}$. Section 11 concludes with directions for further research.
2. Notation
We use the following notation:
$a,b$  $\in\mathbb{C}\text{ (scalar quantified variables)},$
$d,e,f$ $\in\mathbb{N},$
$h$ $\in\{1,\ldots,d\},\ h\neq i,$
$i,j,k$ $\in\{1,\ldots,d\},\{1,\ldots,e\},\{1,\ldots,f\}\text{ respectively},$
$m,n$ $\in\mathbb{N},$
$p,q$ $\in\text{polynomials},$
$r,s$ $\in\mathbb{R},$
$t,u$ $\in\text{terms using }x,$
$v,w$ $\in\mathbb{Q},$
$x$ $\in\mathbb{C}^{n},$
$y,z$ $\in\mathbb{C}.$
The constructions below are uniform in the dimension of $\mathbb{C}^{n}$, and where we need to refer to a vector there, we use $x$. The quantified variables $a,b,r,s,v,w,y,z$ always range over scalars, whether in $\mathbb{C}$, $\mathbb{R}$, or $\mathbb{Q}$.
3. Context
There are three key ingredients in the constructions and proofs here.
First is the algebraic encoding of propositional combinations and inequalities over $\mathbb{R}$, which are used here for $\mathbb{R}$ and as inspiration for $\mathbb{C}$:
Lemma 3.1.
For all $t_{1},t_{2},u\in\mathbb{R}$:
$t_{1}=0\vee t_{2}=0$ $\iff t_{1}t_{2}=0$
$t_{1}=0\wedge t_{2}=0$ $\iff t_{1}^{2}+t_{2}^{2}=0$
$u\neq 0$ $\iff(\exists b)\,bu=1$
$u\geq 0$ $\iff(\exists s)\,u=s^{2}$
$u>0$ $\iff(\exists s)\,s^{2}u=1$
Since only the first and third of these hold over $\mathbb{C}$, an explicit method for encoding propositional structure requires additional ingredients.
The second key ingredient here is Lagrange interpolation. An ordinary example is that, if we want to construct a polynomial $p(a)$ with specified values $t_{1},t_{2},t_{3}$ at the arguments $1,2,3$, we can use
$p(a)=t_{1}\frac{(a-2)(a-3)}{(1-2)(1-3)}+t_{2}\frac{(a-1)(a-3)}{(2-1)(2-3)}+t_{3}\frac{(a-1)(a-2)}{(3-1)(3-2)}.$
The application here is simpler, without the divisions.
Lemma 3.2.
The polynomial
$p(a)=\sum_{i}t_{i}\prod_{h\neq i}(a-h)$
vanishes at $a=i$ iff $t_{i}=0$.
The third key ingredient is basic quantifier elimination over an algebraically closed field, where we need only the elimination of a single existentially quantified equation:
Lemma 3.3.
For any polynomial $q(z)=\sum_{i=1}^{d}q_{i}z^{i}$,
$(\exists z\in\mathbb{C})\,(q(z)=0)$ $\iff q_{0}=0\vee q_{1}\neq 0\vee\cdots\vee q_{d}\neq 0$
$(\forall z\in\mathbb{C})\,(q(z)\neq 0)$ $\iff q_{0}\neq 0\wedge q_{1}=0\wedge\cdots\wedge q_{d}=0$
This restates the fundamental theorem of algebra, and we use it for the impossibility result on $\exists$. We do not use the longer and more complicated results of real quantifier elimination here. (See Swan’s expository article on quantifier elimination over $\mathbb{R}$ and $\mathbb{C}$ for general background, and see Basu’s survey for more details on the complexity, which contrasts with the linear bounds here.)
We call the transformations of this paper Boolean elimination because they provide a dual to quantifier elimination: instead of removing quantifiers at the cost of more propositional complexity, we remove propositional complexity at the cost of a short, fixed quantifier prefix.
4. Normalizing to $\exists\forall$ over $\mathbb{C}$
The constructions in this paper all assume the input formula is presented in conjunctive or disjunctive normal form. The two quantified variables in the resulting $\exists\forall$ and $\forall\exists$ forms correspond to the two Boolean layers in these representations, either disjunctions of conjunctions or conjunctions of disjunctions.
We start here with $\varphi$ in disjunctive normal form:
$\varphi$ $:=\bigvee_{i}\Big(\bigwedge_{j}t_{ij}=0\;\wedge\;\bigwedge_{k}u_{ik}\neq 0\Big)$
The idea is to find a polynomial which encodes each disjunct by a factor that vanishes identically as a polynomial in $b$ iff that disjunct holds.
This $\varphi$ is equivalent to the $\exists\forall$ form below:
$\varphi\iff\exists a\,\forall b\,\prod_{i}\Big[(1-a\prod_{k}u_{ik})+\sum_{j}t_{ij}b^{j}\Big]=0$
For the forward direction, suppose the disjunction holds, and suppose specifically that the $i^{th}$ disjunct holds. We choose $a=\prod_{k}u_{ik}^{-1}$. For that $a$ and for any $b$, the $i^{th}$ factor in the above expression will vanish, so the $\exists\forall$ statement holds.
For the backwards direction, suppose that the $\exists\forall$ statement holds for some $a$. Then for any $b$, one of the factors must vanish, and by the pigeonhole principle, one of the factors must vanish for infinitely many $b$. As a nonzero univariate polynomial has only finitely many roots, if the $i^{th}$ factor vanishes for infinitely many $b$, then it must be the zero polynomial in $b$. Therefore $1-a\prod_{k}u_{ik}=0$, which implies that each $u_{ik}$ is nonzero; and also for each $j$, $t_{ij}=0$. This establishes that the $i^{th}$ disjunct holds, and therefore that the whole disjunction holds.
This proves:
Theorem 4.1.
Let $\varphi(x)$ be a finite combination of polynomial equations and inequations in disjunctive normal form, with $d$ disjuncts and a total of $e$ equations. Then there is a polynomial $p(a,b,x)\in\mathbb{C}[a,b,x]$, of degree $d$ in $a$, of degree $e$ in $b$, and such that for all $x\in\mathbb{C}^{n}$,
$\varphi(x)\Longleftrightarrow(\exists a\in\mathbb{C})(\forall b\in\mathbb{C})\,p(a,b,x)=0.$
5. Normalizing to $\forall\exists$ over $\mathbb{C}$
We start with $\varphi$ in conjunctive normal form:
$\varphi:=\bigwedge_{i}\Big(\bigvee_{j}t_{ij}=0\;\vee\;\bigvee_{k}u_{ik}\neq 0\Big)$
We use $\mathbb{C}$ having characteristic 0 to assign a distinct integer parameter to each conjunct, and apply Lagrange interpolation to construct a polynomial whose $i^{th}$ term contributes only when $a$ equals $i$. Then $\varphi$ is equivalent to the $\forall\exists$ form below:
$\varphi\iff\forall a\,\exists b\,\Big[1-b\prod_{i}(a-i)\Big]\Big[\sum_{i}\prod_{h\neq i}(a-h)\prod_{j}t_{ij}\prod_{k}(1-bu_{ik})\Big]=0.$
For the forward direction, suppose the conjunction holds.
  • If $a$ is distinct from all of the $i$’s, then let $b=\prod_{i}(a-i)^{-1}$, so that the first bracketed expression will vanish.
  • If $a$ is equal to one of the $i$’s, and $t_{ij}=0$, then let $b=0$ (or any other value), and the second bracketed expression will vanish.
  • If $a$ is equal to one of the $i$’s, and $u_{ik}\neq 0$, then let $b=u_{ik}^{-1}$, and the second bracketed expression will vanish.
In any case the $\forall\exists$ statement holds.
For the backwards direction, suppose that the $\forall\exists$ statement holds, and consider $a=i$. Then the big product reduces to
$\prod_{h\neq i}(i-h)\,\prod_{j}t_{ij}\prod_{k}(1-bu_{ik})$
and must vanish. So either $\prod_{j}t_{ij}$ vanishes, in which case one of the $t_{i}$’s vanishes; or $\prod_{k}(1-bu_{ik})$ vanishes, in which case one of the $u_{ik}$ is non-zero. In either case the conjunction holds.
This proves:
Theorem 5.1.
Let $\varphi(x)$ be a finite combination of polynomial equations and inequations in conjunctive normal form, with $d$ conjuncts and at most $f$ inequalities per conjunct. Then there is a polynomial $q(a,b,x)\in\mathbb{C}[a,b,x]$, whose degree in $a$ is 2d12d-1, whose degree in $b$ is at most $f+1$, and such that for all $x\in\mathbb{C}^{n}$,
$\varphi(x)\Longleftrightarrow(\forall a\in\mathbb{C})(\exists b\in\mathbb{C})\,q(a,b,x)=0.$
6. Optimality: the quantifier prefixes cannot be shortened over $\mathbb{C}$
The constructions above used exactly two quantifier blocks. This section shows that such alternation is necessary: no purely universal or purely existential single-equation representation is possible over $\mathbb{C}$. The argument for the existential form is motivated by dimension theory, but we present it in an elementary form that highlights the connection with quasiminimality.
Lemma 6.1.
For any polynomial $p(a,y)$, the set $S=\{y:(\forall a)\,p(a,y)=0\}$ is closed.
Corollary 6.1.
$y\neq 0$ is not equivalent to any formula of the form $(\forall a)\,p(a,y)=0$.
Theorem 6.1.
For any polynomial $p(a,y,z)$, the set
$S=\{(y,z):\exists a,\,p(a,y,z)=0\}$
is either empty or uncountable.
Proof.
Suppose $S$ is countable. We will use the quantifier elimination of $\exists$ over $\mathbb{C}$.
Write
$p(a,y,z)=\sum_{i}p_{i}(y,z)a^{i}.$
For any $(y_{0},z_{0})\notin S$, there is no root for $p(a,y_{0},z_{0})$, so for each $i\geq 1$, $p_{i}$ vanishes outside of $S$. Since $S$ is countable, its complement is dense, and by continuity, $p_{i}$ vanishes on $S$ also. Therefore $p(a,y,z)=p_{0}(y,z)$. Let $q=p_{0}$, and let $S_{y}=\{y:\exists z,\,(y,z)\in S\}$.
Write
$q(y,z)=\sum_{i}q_{i}(y)z^{i}.$
For any $y_{0}\notin S_{y}$, there is no root for $q(y_{0},z)$, so for each $i\geq 1$, $q_{i}$ vanishes outside $S_{y}$. Since $S_{y}$ is countable, its complement is dense, and by continuity, $q_{i}$ vanishes on $S_{y}$ also. Therefore $q(y,z)=q_{0}(y)$.
Therefore $S=\{(y,z):q(y,z)=0\}=\{(y,z):q_{0}(y)=0\}=S_{y}\times\mathbb{C}$. This can only be countable if it is empty. ∎
Corollary 6.2.
$y=0\wedge z=0$ is not equivalent to any formula of the form $(\exists a)\,p(a,y,z)=0$, since $\{(y,z):y=0\wedge z=0\}$ is neither empty nor uncountable.
The impossibility results above are stated for quantification over a single scalar variable, but they extend to blocks of universally or existentially quantified variables. Even in higher dimensions, $S=\{y:(\forall\mathbf{a}\in\mathbb{C}^{m})p(\mathbf{a},y)=0\}$ is closed. Similarly, if
$S=\{(y,z):(\exists\mathbf{a}\in\mathbb{C}^{m})p(\mathbf{a},y,z)=0\}$
then the same argument as above shows that $S$ is either empty or uncountable, by carrying out the argument in each variable one at a time. Consequently, regardless of arity, no normal form with purely existential or purely universal quantification of a single equation can represent all Boolean combinations of polynomial equations over $\mathbb{C}$.
7. Normalizing to $\exists$ over formally real fields
A simpler existential form, with only one new quantifier, is available for any other formally real field, so long as we retain the restriction to formulas with equations and inequations (but not order). This $\exists$ normal form holds for $\mathbb{R}$, $\mathbb{Q}$, and $\mathbb{R}^{\text{alg}}$.
This contrasts with the greater generality of theorems 4.1, 5.1, and 6.1, which work not only over $\mathbb{C}$ but also over any infinite field of characteristic 0. The key properties used there were the invertibility of elements, the infinitude of the field, and the finitude of roots for any nonzero polynomial. Here, by contrast, we use the defining property of a formally real field, namely that a sum of squares is zero iff all the squares are zero.
We switch notation for quantified variables from $a,b$ to $r,s$ to reflect the change from $\mathbb{C}$ to a field like $\mathbb{R}$.
Theorem 7.1.
Let $\varphi(x)$ be a finite combination of polynomial equations and inequations in disjunctive normal form, with $d$ disjuncts. Then there is a polynomial $p(r,x)\in\mathbb{R}[r,x]$, of degree $2d$ in $r$, and such that for all $x\in\mathbb{R}^{n}$,
$\varphi(x)\Longleftrightarrow(\exists r\in\mathbb{R})\,p(r,x)=0.$
Proof.
We start with $\varphi$ in disjunctive normal form:
$\varphi$ $:=\bigvee_{i}\Big(\bigwedge_{j}t_{ij}=0\;\wedge\;\bigwedge_{k}u_{ik}\neq 0\Big)$
This $\varphi$ is equivalent to the $\exists$ form below:
$\varphi\iff\exists r\prod_{i}\Big[\sum_{j}t_{ij}^{2}+(1-r\prod_{k}u_{ik})^{2}\Big]=0$
For the forward direction, suppose the disjunction holds, and specifically that the $i^{th}$ disjunct holds. Then let $r=\prod_{k}u_{ik}^{-1}$, and $\sum_{j}t_{ij}^{2}+(1-r\prod_{k}u_{ik})^{2}=0$, so the existential statement holds.
For the backwards direction, suppose the existential statement holds with some $r$. Then for some $i$, $\sum_{j}t_{ij}^{2}+(1-r\prod_{k}u_{ik})^{2}=0$, which means that each of the $t_{ij}$ vanishes and also $r\prod_{k}u_{ik}=1$, so each of the $u_{ik}$ is non-zero. Therefore the $i^{th}$ disjunct holds and the disjunction holds. ∎
Again with our initial example, now interpreted with $y,z\in\mathbb{R}$,
$(y=0\wedge z\neq 0)\vee(z=0\wedge y\neq 0)\iff\exists r\,\Big[y^{2}+(1-rz)^{2}\Big]\Big[z^{2}+(1-ry)^{2}\Big]=0.$
8. Normalizing to $\exists^{d}$ and $\forall\exists$ over $\mathbb{R}$
For inequalities in $\mathbb{R}$, it is typical to also consider order conditions such as $u>0$. Indeed Boolean combinations polynomial equations and such inequalities over $\mathbb{R}$ can be converted to both $\exists^{d}$ and $\forall\exists$ forms. We state theorems here analogous to the main theorem over $\mathbb{C}$, using the algebraic encoding of order from Lemma 3.1. We omit detailed proofs, since the constructions and proofs are otherwise parallel to the results over $\mathbb{C}$, and we will later cite other results establishing a single-quantifier $\exists$ form for these formulas.
Theorem 8.1.
Let $\varphi(x)$ be a finite combination of polynomial equalities and order-inequalities in conjunctive normal form, with $d$ conjuncts. Then there is a polynomial $p(r_{1},\dots,r_{d},x)\in\mathbb{R}[r_{1},\dots,r_{d},x]$, of degree 4 in each $r_{i}$, and such that, for all $x\in\mathbb{R}^{n}$,
$\varphi(x)\Longleftrightarrow(\exists r_{1},\dots,r_{d}\in\mathbb{R})\,p(r_{1},\dots,r_{d},x)=0.$
Proof Sketch:
$\varphi\ :=\$$\bigwedge_{i}(\bigvee_{j}t_{ij}=0\vee\bigvee_{k}u_{ik}>0)$
[…]
$\iff$ $\exists r_{1},\dots,r_{d}\;\sum_{i}\Big[\prod_{j}t_{ij}\prod_{k}(1-r_{i}^{2}u_{ik})\Big]^{2}=0.$
Theorem 8.2.
Let $\varphi(x)$ be a finite combination of polynomial equalities and order-inequalities in conjunctive normal form, with $d$ conjuncts and at most $f$ inequalities per conjunct. Then there is a polynomial $q(r,s,x)\in\mathbb{R}[r,s,x]$, whose degree in $r$ is 2d12d-1, whose degree in $s$ is at most $2f+1$, and such that for all $x\in\mathbb{R}^{n}$,
$\varphi(x)\Longleftrightarrow(\forall r\in\mathbb{R})(\exists s\in\mathbb{R})\,q(r,s,x)=0.$
Proof Sketch:
$\varphi\ :=\$$\bigwedge_{i}(\bigvee_{j}t_{ij}=0\vee\bigvee_{k}u_{ik}>0)$
$\iff$ $\bigwedge_{i}\exists s\,\prod_{j}t_{ij}\prod_{k}(1-s^{2}u_{ik})=0$
$\iff$ $\forall r\,\exists s\,\Big[1-s\prod_{i}(r-i)\Big]\Big[\sum_{i}\prod_{h\neq i}(r-h)\prod_{j}t_{ij}\prod_{k}(1-s^{2}u_{ik})\Big]=0.$
As an ordered variant of our initial example, if
$\varphi:=(y=0\vee z=0)\wedge(y>0\vee z>0)$
then these constructions yield
$\varphi\iff(\exists r)(\exists s)\Big[yz\Big]^{2}+\Big[(1-s^{2}y)(1-r^{2}z)\Big]^{2}=0$
and
$\varphi\iff(\forall r)(\exists s)\Big[1-s(r-1)(r-2)\Big]\Big[(r-2)yz+(r-1)(1-s^{2}y)(1-s^{2}z)\Big]=0.$
Since there was no way to express inequations such as $u\neq 0$ in a $\forall$ form over $\mathbb{R}$, there is also no way to express inequalities in a $\forall$ form over $\mathbb{R}$.
9. Single-quantifier $\exists$ forms over $\mathbb{R}$
The previous sections gave explicit $\exists^{d}$ and $\forall\exists$ forms for Boolean combinations of inequalities over $\mathbb{R}$. David Marker called our attention to a result of Motzkin showing that in fact a single existential quantifier always suffices. This result was later sharpened by Pecker, who achieved the single quantifier via polynomials with degrees larger than the polynomials above, but much smaller than Motzkin’s. We may write his result in our terms as an $\exists$ normal form:
Theorem 9.1.
Let $\varphi(x)$ be a finite combination of polynomial equalities and $f$ strict inequalities in conjunctive normal form. Then there is a polynomial $p(r,x)\in\mathbb{R}[r,x]$, of degree $2^{f}$ in $r$, such that, for all $x\in\mathbb{R}^{n}$,
$\varphi(x)\Longleftrightarrow(\exists r\in\mathbb{R})\,p(r,x)=0.\quad$
As an example, consider the polynomial produced in Pecker’s first corollary:
$(y>0\wedge z>0)\iff(\exists r)(r^{2}yz-y^{2}z-1)^{2}-y^{2}z=0$
We may simplify this to
$(y>0\wedge z>0)\iff(\exists r)(r^{2}yz-z-1)^{2}=z$
Solving the simplified equation for $y$ gives $y=(z+1\pm\sqrt{z})/r^{2}z$, which makes it easier to see that it has roots exactly when $y$ and $z$ are positive.
As $r$ takes all possible values, graphs like the below sweep through the entire first quadrant.
Figure 1. Graph for a simple version of Pecker’s polynomial
A contrast between Pecker’s results and ours is that our results, by using two new variables, require only one use of each term $t_{ij}$ or $u_{ik}$, and have linearly growing degrees. Pecker’s results, using only one new variable, require multiple uses of those terms, and have exponentially growing degrees.
10. Normal forms over $\mathbb{Q}$
The results over $\mathbb{R}$ hold with some modifications over $\mathbb{Q}$, since not every positive rational is a square. Analogous results obtain from replacing the squares in $\mathbb{R}$ with sums of four squares in $\mathbb{Q}$. Thus Pecker’s form yields an $\exists^{4}$ normal form for $\mathbb{Q}$, and the forms provided explicitly above yield $\exists^{4d}$ and $\forall\exists^{4}$ forms over $\mathbb{Q}$.
We can also reduce the $4$’s to $3$’s in these signatures, by using a fact from number theory and a corollary reformulation. We use $v,w$ for quantified variables to reflect the switch from $\mathbb{C}$ and $\mathbb{R}$ to $\mathbb{Q}$:
Lemma 10.1.
For any natural number $n$, either $n$ or $2n$ is the sum of three squares.
Proof.
If $n$ is not of the form $4^{k}(8m+7)$, then by Legendre’s theorem $n$ is a sum of three squares. If $n$ is $4^{k}(8m+7)$, then $2n$ is not of that form, so $2n$ is a sum of three squares. ∎
Corollary 10.1.
$u\in\mathbb{Q}\wedge u>0\iff(\exists v_{1},v_{2},v_{3}\in\mathbb{Q})\,[1-u(v_{1}^{2}+v_{2}^{2}+v_{3}^{2})][1-2u(v_{1}^{2}+v_{2}^{2}+v_{3}^{2})]=0\Big$.
Proof.
The left-to-right direction follows from the lemma with $n$ as the product of the numerator and denominator of $1/(2u)$. The right-to-left direction holds because any sum of three squares is non-negative, and twice such a sum is non-negative also. ∎
In particular, this leads to the following $\exists^{3d}$ and $\forall\exists^{3}$ normal forms over $\mathbb{Q}$: For $\varphi$ in conjunctive normal form as above, we have
$\varphi\iff\exists v_{1},\dots,v_{3d}\;\sum_{i}\Big[\prod_{j}t_{ij}\prod_{k}(1-u_{ik}V_{i})(1-2u_{ik}V_{i})\Big]^{2}=0.$
where $V_{i}=v_{3i-2}^{2}+v_{3i-1}^{2}+v_{3i}^{2}$. Also
$\varphi\iff\forall v\,\exists w_{1},w_{2},w_{3}\,$ $\Big[1-w_{1}\prod_{i}(v-i)\Big]\Big[\sum_{i}\prod_{h\neq i}(v-h)\prod_{j}t_{ij}\prod_{k}(1-u_{ik}W)(1-2u_{ik}W)\Big]=0.$
where $W=w_{1}^{2}+w_{2}^{2}+w_{3}^{2}$.
These forms may or may not be minimal; we provide them to show the possibility of a number-theoretic Boolean elimination over $\mathbb{Q}$, where quantifier elimination is not available.
11. Conclusion
The results here show that the Boolean structure of constructible or semialgebraic sets can be uniformly represented with a single polynomial equation under either an $\exists\forall$ or a $\forall\exists$ prefix for $\mathbb{C}$, or even under a simple $\exists$ prefix for $\mathbb{R}$, and that those quantifier patterns are optimal. We hope it will be useful to have those results conveniently assembled here.
These results also lead to questions about various extensions of them.
  • Can the $\exists$ results about $\mathbb{R}$ be simplified or optimized for smaller degrees, and can the $\exists^{3}$ results about $\mathbb{Q}$ be reduced to fewer quantifiers?
  • Can every universally quantified formula be normalized to $\forall\exists$ form with a single equation, and can every existentially quantified formula be normalized to $\exists\forall$ form?
  • What variants of these constructions are available over $\mathbb{C}$ when the Boolean combinations may be countable as in $\mathcal{L}_{\omega_{1},\omega}$ rather than finite, and the language may be expanded beyond polynomials to include exponentials or other univariate entire functions?
These questions point beyond elementary methods toward a deeper study of Boolean elimination.